A train is moving towards north and a bird is flying towards south above train in opposite direction. Find length of the train if parrot crossed train in 2min. Velocity of train is 108km/h and velocity of bird is 18km/h.

Grade:11

1 Answers

Eshan
5 years ago
Dear student,

The relative velocity of the bird with respect to train is 18km/h-(-108km/hr)=126km/hr
Hence from the frame of reference the train, it observes that the bird cross all of its length in 2 minutes with a speed of 126km/hr.

Hence the length of the train isv\times t=126km/hr\times \dfrac{2}{60}hr=4.2km